OpenStudy (anonymous):

who can help me with math i have 3 questions left and have to solve in the dark so please help the best you can

OpenStudy (anonymous):

@Teddyiswatshecallsme

OpenStudy (anonymous):

Post your questions.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Next questions?

OpenStudy (anonymous):

midpoint formula : (x1 + x2) / 2 , (y1 + y2)/2 (3,3)....x1 = 3 and y1 = 3 (1,-1)..x2 = 1 and y2 = -1 now we sub (3 + 1) / 2 , (3 + (-1) / 2 (4/2) , (2/2) (2,1) is the midpoint ....yep, its C

OpenStudy (anonymous):

thx guys thts my nxt 1

OpenStudy (sleepyjess):

distance formula is \(\sqrt{(x_2-x_1)^2+(y_2-y_1)}\) 6, 4 = \(x_1,~y_1\) 2, 3 = \(x_2,~y_2\)

OpenStudy (sleepyjess):

substitute in the values and you have \(\sqrt{2-6)^2+(3-4)^2}\)

OpenStudy (anonymous):

Oh! So it's D

OpenStudy (anonymous):

sqrt of 7?

OpenStudy (sleepyjess):

whoops should be (2-6)^2

OpenStudy (sleepyjess):

2-6=-4 -4*-4=16 3-4=-1 -1*-1=1 16+1=17 \(\sqrt{17}\)

OpenStudy (anonymous):

oh yeah 17 sorry guysim messing up cuz im in the dark

OpenStudy (anonymous):

last question tho.....

OpenStudy (anonymous):

Mhmm. . .

OpenStudy (anonymous):

C

OpenStudy (sleepyjess):

midpoint = \(\dfrac{x_1~+~x_2}{2},~\dfrac{y_1~+~y_2}{2}~6,~4~=~x_1,~y_1~~~2,~3~=~x_2,~y_2\)

OpenStudy (anonymous):

thanks Ethereal and sleepyjess :)

OpenStudy (anonymous):

Yes, especially Jess. ^^

OpenStudy (sleepyjess):

\(\dfrac{6+2}{2},~\dfrac{4+3}{2}=\dfrac{8}{2},~\dfrac{7}{2}=4,~\dfrac{7}{2}\)